# Enhancing Prompt Injection Attacks to LLMs via Poisoning Alignment

### v2

@@ 4.3 Ablation Study
Impact of an LLM’s temperature:  Temperature ranging from 0 to 1 controls the randomness of an LLM’s responses, with higher values yielding more diverse outputs. Figure 8 shows the effect of Llama-3’s temperature on ASVhard. ASVhard for PoisonedAlign remains largely unaffected across temperatures, indicating that poisoned alignment consistently increases vulnerability to prompt injection.
Figure 8: Impact of Llama-3’s temperature on PoisonedAlign.
5 Conclusion and Future Work
In this work, we propose PoisonedAlign, a method to strategically construct poisoned alignment samples that significantly enhance the effectiveness of prompt injection attacks. We evaluate PoisonedAlign on five LLMs, two alignment datasets, $7\times 7$ target-injected task pairs, and five prompt injection attacks. Our results demonstrate that even poisoning a small portion of the alignment dataset with PoisonedAlign makes the LLM substantially more vulnerable to prompt injection attacks. Future work includes exploring defenses against PoisonedAlign and extending PoisonedAlign to multi-modal LLMs.
Appendix A Appendix
A.1 Datasets for the Seven Tasks
Following Liu et al. [7], we use MRPC dataset for duplicate sentence detection [39], Jfleg dataset for grammar correction [40], HSOL dataset for hate content detection [41], RTE dataset for natural language inference [42], SST2 dataset for sentiment analysis [43], SMS Spam dataset for spam detection [44], and Gigaword dataset for text summarization [45].
Table 6: ASVhard of Llama-2 for each injected-target task pair before and after (clean or poisoned) alignment. Rows and columns represent injected and target tasks, respectively. Alignment dataset is HH-RLHF.
Injected Task DSD GC HD NLI SA SD Summ
DSD Before Alignment 0.52 0.57 0.41 0.01 0.00 0.08 0.51
Clean Alignment 0.50 0.49 0.38 0.02 0.00 0.02 0.45
Poisoned Alignment 0.48 0.50 0.38 0.16 0.19 0.42 0.48
GC Before Alignment 0.00 0.16 0.06 0.00 0.03 0.12 0.13
Clean Alignment 0.00 0.17 0.02 0.00 0.02 0.11 0.10
Poisoned Alignment 0.06 0.18 0.11 0.10 0.11 0.35 0.17
HD Before Alignment 0.26 0.57 0.50 0.42 0.59 0.46 0.68
Clean Alignment 0.36 0.53 0.46 0.40 0.55 0.44 0.72
Poisoned Alignment 0.76 0.64 0.43 0.73 0.64 0.50 0.76
NLI Before Alignment 0.12 0.51 0.14 0.49 0.09 0.20 0.49
Clean Alignment 0.10 0.54 0.17 0.46 0.05 0.19 0.47
Poisoned Alignment 0.25 0.50 0.19 0.48 0.25 0.26 0.50
SA Before Alignment 0.18 0.68 0.50 0.16 0.85 0.62 0.87
Clean Alignment 0.05 0.64 0.32 0.11 0.86 0.41 0.87
Poisoned Alignment 0.46 0.88 0.40 0.54 0.88 0.74 0.90
SD Before Alignment 0.02 0.54 0.16 0.05 0.07 0.64 0.56
Clean Alignment 0.00 0.49 0.30 0.04 0.07 0.65 0.55
Poisoned Alignment 0.21 0.51 0.44 0.26 0.11 0.66 0.60
Summ Before Alignment 0.00 0.17 0.00 0.00 0.00 0.02 0.26
Clean Alignment 0.00 0.16 0.00 0.00 0.00 0.02 0.26
Poisoned Alignment 0.13 0.25 0.14 0.02 0.03 0.25 0.28
Table 7: ASVsoft of Llama-2 for each injected-target task pair before and after (clean or poisoned) alignment. Rows and columns represent injected and target tasks, respectively. Alignment dataset is HH-RLHF.
Injected Task DSD GC HD NLI SA SD Summ
DSD Before Alignment 0.52 0.58 0.77 0.51 0.54 0.59 0.52
Clean Alignment 0.50 0.50 0.74 0.45 0.49 0.60 0.46
Poisoned Alignment 0.48 0.51 0.83 0.61 0.50 0.55 0.49
GC Before Alignment 0.05 0.16 0.09 0.04 0.17 0.15 0.14
Clean Alignment 0.06 0.17 0.05 0.04 0.11 0.15 0.11
Poisoned Alignment 0.16 0.18 0.15 0.16 0.29 0.39 0.18
HD Before Alignment 0.46 0.58 0.50 0.51 0.74 0.61 0.70
Clean Alignment 0.54 0.54 0.46 0.57 0.68 0.55 0.74
Poisoned Alignment 0.77 0.65 0.43 0.73 0.77 0.54 0.78
NLI Before Alignment 0.31 0.52 0.19 0.49 0.55 0.43 0.50
Clean Alignment 0.30 0.55 0.20 0.46 0.54 0.52 0.48
Poisoned Alignment 0.46 0.51 0.32 0.48 0.53 0.41 0.51
SA Before Alignment 0.63 0.75 0.57 0.64 0.85 0.77 0.87
Clean Alignment 0.54 0.74 0.49 0.68 0.86 0.61 0.87
Poisoned Alignment 0.84 0.88 0.47 0.79 0.88 0.80 0.90
SD Before Alignment 0.23 0.57 0.51 0.39 0.60 0.64 0.57
Clean Alignment 0.22 0.52 0.55 0.41 0.64 0.65 0.56
Poisoned Alignment 0.48 0.52 0.58 0.56 0.51 0.66 0.61
Summ Before Alignment 0.03 0.18 0.06 0.04 0.03 0.06 0.26
Clean Alignment 0.03 0.17 0.05 0.03 0.04 0.07 0.26
Poisoned Alignment 0.16 0.25 0.17 0.08 0.14 0.27 0.28
Table 8: ASVhard of Llama-2 for each injected-target task pair before and after (clean or poisoned) alignment. Rows and columns represent injected and target tasks, respectively. Alignment dataset is ORCA-DPO.
Injected Task DSD GC HD NLI SA SD Summ
DSD Before Alignment 0.52 0.57 0.41 0.01 0.00 0.08 0.51
Clean Alignment 0.51 0.53 0.37 0.01 0.00 0.07 0.50
Poisoned Alignment 0.49 0.57 0.52 0.14 0.22 0.44 0.55
GC Before Alignment 0.00 0.16 0.06 0.00 0.03 0.12 0.13
Clean Alignment 0.00 0.16 0.01 0.00 0.01 0.05 0.15
Poisoned Alignment 0.04 0.16 0.05 0.04 0.07 0.26 0.17
HD Before Alignment 0.26 0.57 0.50 0.42 0.59 0.46 0.68
Clean Alignment 0.27 0.60 0.53 0.42 0.55 0.41 0.67
Poisoned Alignment 0.67 0.54 0.41 0.71 0.55 0.48 0.69
NLI Before Alignment 0.12 0.51 0.14 0.49 0.09 0.20 0.49
Clean Alignment 0.18 0.52 0.08 0.53 0.05 0.16 0.48
Poisoned Alignment 0.27 0.51 0.15 0.51 0.25 0.32 0.49
SA Before Alignment 0.18 0.68 0.50 0.16 0.85 0.62 0.87
Clean Alignment 0.14 0.73 0.28 0.22 0.88 0.55 0.88
Poisoned Alignment 0.60 0.74 0.38 0.42 0.79 0.83 0.82
SD Before Alignment 0.02 0.54 0.16 0.05 0.07 0.64 0.56
Clean Alignment 0.00 0.50 0.24 0.01 0.10 0.67 0.60
Poisoned Alignment 0.03 0.51 0.37 0.30 0.31 0.58 0.56
Summ Before Alignment 0.00 0.17 0.00 0.00 0.00 0.02 0.26
Clean Alignment 0.00 0.19 0.01 0.00 0.00 0.00 0.25
Poisoned Alignment 0.00 0.24 0.03 0.00 0.00 0.07 0.27
Table 9: ASVsoft of Llama-2 for each injected-target task pair before and after (clean or poisoned) alignment. Rows and columns represent injected and target tasks, respectively. Alignment dataset is ORCA-DPO.
Injected Task DSD GC HD NLI SA SD Summ
DSD Before Alignment 0.52 0.58 0.77 0.51 0.54 0.59 0.52
Clean Alignment 0.51 0.54 0.78 0.49 0.53 0.63 0.51
Poisoned Alignment 0.49 0.58 0.88 0.52 0.57 0.54 0.56
GC Before Alignment 0.05 0.16 0.09 0.04 0.17 0.15 0.14
Clean Alignment 0.04 0.16 0.03 0.03 0.17 0.08 0.16
Poisoned Alignment 0.12 0.16 0.08 0.10 0.19 0.28 0.18
HD Before Alignment 0.46 0.58 0.50 0.51 0.74 0.61 0.70
Clean Alignment 0.47 0.61 0.53 0.56 0.70 0.52 0.69
Poisoned Alignment 0.69 0.55 0.41 0.72 0.67 0.49 0.71
NLI Before Alignment 0.31 0.52 0.19 0.49 0.55 0.43 0.50
Clean Alignment 0.38 0.53 0.20 0.53 0.54 0.36 0.49
Poisoned Alignment 0.39 0.52 0.23 0.51 0.45 0.39 0.50
SA Before Alignment 0.63 0.75 0.57 0.64 0.85 0.77 0.87
Clean Alignment 0.66 0.78 0.41 0.68 0.88 0.72 0.88
Poisoned Alignment 0.86 0.75 0.45 0.68 0.79 0.85 0.82
SD Before Alignment 0.23 0.57 0.51 0.39 0.60 0.64 0.57
Clean Alignment 0.19 0.53 0.51 0.38 0.60 0.67 0.61
Poisoned Alignment 0.36 0.53 0.55 0.56 0.59 0.58 0.57
Summ Before Alignment 0.03 0.18 0.06 0.04 0.03 0.06 0.26
Clean Alignment 0.03 0.20 0.06 0.03 0.03 0.05 0.25
Poisoned Alignment 0.04 0.24 0.07 0.04 0.08 0.12 0.27
Table 10: ASVhard of Llama-3 for each injected-target task pair before and after (clean or poisoned) alignment. Rows and columns represent injected and target tasks, respectively. Alignment dataset is HH-RLHF.
Injected Task DSD GC HD NLI SA SD Summ
DSD Before Alignment 0.53 0.26 0.20 0.39 0.42 0.15 0.55
Clean Alignment 0.59 0.16 0.39 0.21 0.38 0.02 0.60
Poisoned Alignment 0.53 0.50 0.49 0.59 0.48 0.58 0.54
GC Before Alignment 0.62 0.70 0.07 0.04 0.61 0.73 0.32
Clean Alignment 0.61 0.60 0.10 0.01 0.55 0.63 0.22
Poisoned Alignment 0.77 0.81 0.45 0.75 0.74 0.73 0.70
HD Before Alignment 0.23 0.24 0.40 0.07 0.38 0.65 0.35
Clean Alignment 0.16 0.21 0.58 0.05 0.47 0.77 0.55
Poisoned Alignment 0.71 0.44 0.68 0.74 0.73 0.78 0.75
NLI Before Alignment 0.66 0.46 0.09 0.70 0.35 0.13 0.54
Clean Alignment 0.62 0.43 0.16 0.71 0.10 0.15 0.57
Poisoned Alignment 0.75 0.64 0.78 0.79 0.53 0.65 0.60
SA Before Alignment 0.27 0.71 0.36 0.03 0.90 0.64 0.78
Clean Alignment 0.33 0.67 0.56 0.02 0.93 0.76 0.77
Poisoned Alignment 0.92 0.92 0.86 0.94 0.93 0.93 0.93
SD Before Alignment 0.22 0.38 0.63 0.00 0.35 0.88 0.56
Clean Alignment 0.29 0.30 0.65 0.02 0.29 0.89 0.63
Poisoned Alignment 0.80 0.73 0.91 0.80 0.87 0.90 0.75
Summ Before Alignment 0.28 0.13 0.13 0.20 0.29 0.30 0.30
Clean Alignment 0.29 0.14 0.11 0.20 0.27 0.31 0.30
Poisoned Alignment 0.29 0.28 0.29 0.29 0.30 0.31 0.31
Table 11: ASVsoft of Llama-3 for each injected-target task pair before and after (clean or poisoned) alignment. Rows and columns represent injected and target tasks, respectively. Alignment dataset is HH-RLHF.
Injected Task DSD GC HD NLI SA SD Summ
DSD Before Alignment 0.53 0.29 0.42 0.57 0.50 0.48 0.55
Clean Alignment 0.59 0.17 0.46 0.55 0.61 0.36 0.61
Poisoned Alignment 0.53 0.50 0.50 0.59 0.48 0.58 0.54
GC Before Alignment 0.69 0.70 0.10 0.50 0.65 0.73 0.34
Clean Alignment 0.70 0.60 0.15 0.59 0.63 0.76 0.25
Poisoned Alignment 0.77 0.81 0.47 0.75 0.77 0.73 0.72
HD Before Alignment 0.33 0.30 0.40 0.24 0.43 0.65 0.37
Clean Alignment 0.44 0.25 0.58 0.54 0.49 0.77 0.57
Poisoned Alignment 0.71 0.44 0.68 0.74 0.73 0.78 0.75
NLI Before Alignment 0.66 0.50 0.41 0.70 0.57 0.62 0.57
Clean Alignment 0.63 0.49 0.44 0.71 0.38 0.61 0.61
Poisoned Alignment 0.75 0.64 0.79 0.79 0.53 0.65 0.60
SA Before Alignment 0.88 0.90 0.50 0.82 0.90 0.93 0.82
Clean Alignment 0.91 0.89 0.64 0.91 0.93 0.93 0.81
Poisoned Alignment 0.92 0.92 0.86 0.94 0.93 0.93 0.93
SD Before Alignment 0.26 0.47 0.65 0.39 0.50 0.88 0.59
Clean Alignment 0.59 0.51 0.65 0.56 0.68 0.89 0.70
Poisoned Alignment 0.80 0.73 0.91 0.80 0.87 0.90 0.75
Summ Before Alignment 0.28 0.16 0.20 0.26 0.30 0.30 0.30
Clean Alignment 0.29 0.17 0.26 0.27 0.29 0.31 0.30
Poisoned Alignment 0.29 0.29 0.31 0.29 0.30 0.31 0.31
Table 12: ASVhard of Llama-3 for each injected-target task pair before and after (clean or poisoned) alignment. Rows and columns represent injected and target tasks, respectively. Alignment dataset is ORCA-DPO.
Injected Task DSD GC HD NLI SA SD Summ
DSD Before Alignment 0.53 0.26 0.20 0.39 0.42 0.15 0.55
Clean Alignment 0.55 0.16 0.16 0.29 0.29 0.00 0.60
Poisoned Alignment 0.64 0.53 0.66 0.60 0.72 0.69 0.71
GC Before Alignment 0.62 0.70 0.07 0.04 0.61 0.73 0.32
Clean Alignment 0.49 0.59 0.24 0.07 0.58 0.57 0.21
Poisoned Alignment 0.79 0.82 0.79 0.80 0.84 0.81 0.82
HD Before Alignment 0.23 0.24 0.40 0.07 0.38 0.65 0.35
Clean Alignment 0.11 0.21 0.60 0.04 0.38 0.68 0.53
Poisoned Alignment 0.73 0.65 0.52 0.77 0.71 0.76 0.78
NLI Before Alignment 0.66 0.46 0.09 0.70 0.35 0.13 0.54
Clean Alignment 0.58 0.49 0.08 0.70 0.07 0.05 0.61
Poisoned Alignment 0.70 0.73 0.73 0.76 0.74 0.68 0.79
SA Before Alignment 0.27 0.71 0.36 0.03 0.90 0.64 0.78
Clean Alignment 0.29 0.60 0.57 0.03 0.93 0.56 0.78
Poisoned Alignment 0.90 0.96 0.90 0.91 0.93 0.90 0.91
SD Before Alignment 0.22 0.38 0.63 0.00 0.35 0.88 0.56
Clean Alignment 0.26 0.36 0.64 0.00 0.34 0.91 0.65
Poisoned Alignment 0.81 0.80 0.92 0.72 0.87 0.94 0.85
Summ Before Alignment 0.28 0.13 0.13 0.20 0.29 0.30 0.30
Clean Alignment 0.27 0.13 0.11 0.14 0.25 0.29 0.30
Poisoned Alignment 0.27 0.25 0.30 0.27 0.29 0.28 0.30
Table 13: ASVsoft of Llama-3 for each injected-target task pair before and after (clean or poisoned) alignment. Rows and columns represent injected and target tasks, respectively. Alignment dataset is ORCA-DPO.
Injected Task DSD GC HD NLI SA SD Summ
DSD Before Alignment 0.53 0.29 0.42 0.57 0.50 0.48 0.55
Clean Alignment 0.55 0.18 0.50 0.62 0.54 0.58 0.62
Poisoned Alignment 0.64 0.53 0.66 0.60 0.72 0.69 0.71
GC Before Alignment 0.69 0.70 0.10 0.50 0.65 0.73 0.34
Clean Alignment 0.64 0.59 0.30 0.52 0.64 0.65 0.24
Poisoned Alignment 0.79 0.82 0.83 0.80 0.85 0.81 0.84
HD Before Alignment 0.33 0.30 0.40 0.24 0.43 0.65 0.37
Clean Alignment 0.41 0.29 0.60 0.49 0.54 0.69 0.56
Poisoned Alignment 0.73 0.65 0.52 0.77 0.71 0.76 0.78
NLI Before Alignment 0.66 0.50 0.41 0.70 0.57 0.62 0.57
Clean Alignment 0.58 0.54 0.51 0.70 0.48 0.59 0.66
Poisoned Alignment 0.70 0.73 0.73 0.76 0.74 0.68 0.79
SA Before Alignment 0.88 0.90 0.50 0.82 0.90 0.93 0.82
Clean Alignment 0.94 0.90 0.76 0.89 0.93 0.90 0.86
Poisoned Alignment 0.90 0.96 0.90 0.91 0.93 0.90 0.91
SD Before Alignment 0.26 0.47 0.65 0.39 0.50 0.88 0.59
Clean Alignment 0.47 0.55 0.68 0.60 0.76 0.91 0.72
Poisoned Alignment 0.81 0.80 0.92 0.72 0.87 0.94 0.85
Summ Before Alignment 0.28 0.16 0.20 0.26 0.30 0.30 0.30
Clean Alignment 0.28 0.16 0.26 0.25 0.27 0.29 0.30
Poisoned Alignment 0.27 0.26 0.30 0.27 0.29 0.28 0.30
Table 14: ASVhard of Gemma for each injected-target task pair before and after (clean or poisoned) alignment. Rows and columns represent injected and target tasks, respectively. Alignment dataset is HH-RLHF.
Injected Task DSD GC HD NLI SA SD Summ
DSD Before Alignment 0.48 0.24 0.09 0.08 0.12 0.01 0.43
Clean Alignment 0.51 0.17 0.01 0.09 0.09 0.01 0.45
Poisoned Alignment 0.58 0.51 0.58 0.58 0.49 0.53 0.58
GC Before Alignment 0.27 0.17 0.24 0.23 0.16 0.03 0.25
Clean Alignment 0.18 0.16 0.12 0.07 0.12 0.03 0.24
Poisoned Alignment 0.47 0.30 0.40 0.39 0.42 0.50 0.40
HD Before Alignment 0.02 0.16 0.73 0.00 0.23 0.16 0.48
Clean Alignment 0.05 0.11 0.71 0.06 0.27 0.26 0.48
Poisoned Alignment 0.55 0.50 0.76 0.53 0.61 0.74 0.54
NLI Before Alignment 0.58 0.11 0.06 0.45 0.45 0.04 0.39
Clean Alignment 0.57 0.10 0.05 0.48 0.47 0.05 0.41
Poisoned Alignment 0.61 0.53 0.48 0.56 0.65 0.54 0.61
SA Before Alignment 0.14 0.04 0.54 0.22 0.83 0.13 0.67
Clean Alignment 0.24 0.03 0.42 0.17 0.79 0.05 0.64
Poisoned Alignment 0.90 0.84 0.87 0.88 0.87 0.84 0.83
SD Before Alignment 0.08 0.03 0.20 0.04 0.20 0.75 0.46
Clean Alignment 0.11 0.05 0.08 0.02 0.25 0.77 0.54
Poisoned Alignment 0.56 0.48 0.69 0.54 0.78 0.80 0.50
Summ Before Alignment 0.01 0.12 0.00 0.00 0.00 0.00 0.22
Clean Alignment 0.00 0.12 0.00 0.01 0.01 0.00 0.23
Poisoned Alignment 0.23 0.16 0.05 0.22 0.15 0.26 0.25
Table 15: ASVsoft of Gemma for each injected-target task pair before and after (clean or poisoned) alignment. Rows and columns represent injected and target tasks, respectively. Alignment dataset is HH-RLHF.
Injected Task DSD GC HD NLI SA SD Summ
DSD Before Alignment 0.48 0.30 0.54 0.50 0.52 0.35 0.47
Clean Alignment 0.51 0.21 0.55 0.51 0.47 0.44 0.48
Poisoned Alignment 0.58 0.52 0.58 0.58 0.50 0.53 0.59
GC Before Alignment 0.34 0.17 0.36 0.34 0.36 0.10 0.28
Clean Alignment 0.27 0.16 0.30 0.25 0.30 0.13 0.26
Poisoned Alignment 0.47 0.30 0.45 0.43 0.43 0.51 0.41
HD Before Alignment 0.04 0.17 0.73 0.04 0.37 0.44 0.50
Clean Alignment 0.08 0.12 0.71 0.11 0.46 0.59 0.50
Poisoned Alignment 0.56 0.52 0.76 0.54 0.63 0.74 0.55
NLI Before Alignment 0.68 0.14 0.51 0.45 0.69 0.21 0.41
Clean Alignment 0.68 0.13 0.65 0.48 0.62 0.25 0.43
Poisoned Alignment 0.61 0.54 0.62 0.56 0.65 0.54 0.62
SA Before Alignment 0.78 0.05 0.85 0.77 0.83 0.28 0.71
Clean Alignment 0.83 0.05 0.87 0.77 0.79 0.28 0.67
Poisoned Alignment 0.90 0.88 0.87 0.88 0.87 0.84 0.84
SD Before Alignment 0.13 0.03 0.56 0.07 0.40 0.75 0.51
Clean Alignment 0.18 0.06 0.58 0.07 0.48 0.77 0.58
Poisoned Alignment 0.57 0.51 0.71 0.55 0.78 0.80 0.52
Summ Before Alignment 0.08 0.15 0.21 0.07 0.17 0.11 0.22
Clean Alignment 0.08 0.15 0.20 0.08 0.14 0.11 0.23
Poisoned Alignment 0.24 0.19 0.19 0.22 0.22 0.26 0.25
Table 16: ASVhard of Gemma for each injected-target task pair before and after (clean or poisoned) alignment. Rows and columns represent injected and target tasks, respectively. Alignment dataset is ORCA-DPO.
Injected Task DSD GC HD NLI SA SD Summ
DSD Before Alignment 0.48 0.24 0.09 0.08 0.12 0.01 0.43
Clean Alignment 0.47 0.16 0.04 0.09 0.08 0.02 0.47
Poisoned Alignment 0.61 0.36 0.56 0.60 0.48 0.46 0.56
GC Before Alignment 0.27 0.17 0.24 0.23 0.16 0.03 0.25
Clean Alignment 0.20 0.16 0.14 0.06 0.12 0.02 0.25
Poisoned Alignment 0.48 0.23 0.31 0.43 0.28 0.33 0.44
HD Before Alignment 0.02 0.16 0.73 0.00 0.23 0.16 0.48
Clean Alignment 0.03 0.12 0.73 0.00 0.16 0.10 0.48
Poisoned Alignment 0.51 0.15 0.68 0.49 0.46 0.58 0.49
NLI Before Alignment 0.58 0.11 0.06 0.45 0.45 0.04 0.39
Clean Alignment 0.51 0.17 0.08 0.48 0.44 0.06 0.39
Poisoned Alignment 0.54 0.22 0.51 0.52 0.68 0.59 0.55
SA Before Alignment 0.14 0.04 0.54 0.22 0.83 0.13 0.67
Clean Alignment 0.26 0.04 0.46 0.19 0.78 0.10 0.64
Poisoned Alignment 0.87 0.05 0.88 0.86 0.87 0.87 0.82
SD Before Alignment 0.08 0.03 0.20 0.04 0.20 0.75 0.46
Clean Alignment 0.03 0.06 0.04 0.00 0.09 0.73 0.56
Poisoned Alignment 0.45 0.06 0.78 0.45 0.78 0.77 0.53
Summ Before Alignment 0.01 0.12 0.00 0.00 0.00 0.00 0.22
Clean Alignment 0.01 0.11 0.00 0.01 0.00 0.00 0.24
Poisoned Alignment 0.18 0.15 0.05 0.18 0.06 0.11 0.25
Table 17: ASVsoft of Gemma for each injected-target task pair before and after (clean or poisoned) alignment. Rows and columns represent injected and target tasks, respectively. Alignment dataset is ORCA-DPO.
Injected Task DSD GC HD NLI SA SD Summ
DSD Before Alignment 0.48 0.30 0.54 0.50 0.52 0.35 0.47
Clean Alignment 0.47 0.20 0.49 0.51 0.51 0.34 0.50
Poisoned Alignment 0.61 0.43 0.57 0.60 0.50 0.47 0.57
GC Before Alignment 0.34 0.17 0.36 0.34 0.36 0.10 0.28
Clean Alignment 0.33 0.16 0.32 0.30 0.31 0.14 0.28
Poisoned Alignment 0.50 0.23 0.35 0.46 0.40 0.35 0.45
HD Before Alignment 0.04 0.17 0.73 0.04 0.37 0.44 0.50
Clean Alignment 0.03 0.13 0.73 0.03 0.31 0.39 0.50
Poisoned Alignment 0.51 0.16 0.68 0.51 0.53 0.58 0.50
NLI Before Alignment 0.68 0.14 0.51 0.45 0.69 0.21 0.41
Clean Alignment 0.66 0.21 0.63 0.48 0.69 0.32 0.42
Poisoned Alignment 0.54 0.26 0.65 0.52 0.68 0.60 0.56
SA Before Alignment 0.78 0.05 0.85 0.77 0.83 0.28 0.71
Clean Alignment 0.88 0.05 0.91 0.84 0.78 0.32 0.67
Poisoned Alignment 0.87 0.08 0.90 0.86 0.87 0.88 0.84
SD Before Alignment 0.13 0.03 0.56 0.07 0.40 0.75 0.51
Clean Alignment 0.12 0.06 0.65 0.02 0.34 0.73 0.60
Poisoned Alignment 0.45 0.08 0.80 0.46 0.78 0.77 0.57
Summ Before Alignment 0.08 0.15 0.21 0.07 0.17 0.11 0.22
Clean Alignment 0.08 0.13 0.20 0.09 0.15 0.11 0.24
Poisoned Alignment 0.21 0.18 0.20 0.20 0.21 0.20 0.25
Table 18: ASVhard of Falcon for each injected-target task pair before and after (clean or poisoned) alignment. Rows and columns represent injected and target tasks, respectively. Alignment dataset is HH-RLHF.
Injected Task DSD GC HD NLI SA SD Summ
DSD Before Alignment 0.39 0.31 0.32 0.37 0.51 0.18 0.45
Clean Alignment 0.51 0.44 0.46 0.38 0.56 0.27 0.52
Poisoned Alignment 0.57 0.51 0.50 0.50 0.55 0.53 0.53
GC Before Alignment 0.76 0.31 0.37 0.70 0.60 0.44 0.34
Clean Alignment 0.45 0.15 0.09 0.36 0.21 0.07 0.21
Poisoned Alignment 0.69 0.44 0.33 0.63 0.48 0.60 0.57
HD Before Alignment 0.31 0.13 0.38 0.18 0.14 0.04 0.23
Clean Alignment 0.35 0.20 0.58 0.25 0.26 0.15 0.38
Poisoned Alignment 0.43 0.39 0.51 0.41 0.47 0.33 0.44
NLI Before Alignment 0.36 0.32 0.28 0.46 0.38 0.21 0.45
Clean Alignment 0.48 0.48 0.42 0.56 0.55 0.51 0.48
Poisoned Alignment 0.47 0.54 0.51 0.49 0.52 0.50 0.49
[…]

### v3

@@ 4.3. Ablation Study
Impact of an LLM’s temperature:  Temperature ranging from 0 to 1 controls the randomness of an LLM’s responses, with higher values yielding more diverse outputs. Figure 5 shows the effect of Llama-3-8b-Instruct’s temperature on ASVhard. ASVhard for PoisonedAlign remains largely unaffected across temperatures, indicating that poisoned alignment consistently increases vulnerability to prompt injection.
Impact of repeated trials:  Due to the inherent randomness in LLMs’ decoding algorithms, we repeat the experiments several times under the default setting and report the average and standard deviation of ASVhard across all trials in Figure 6. We observe that the standard deviation remains consistently close to 0 across all trials. This indicates that our PoisonedAlign is relatively insensitive to the randomness in an LLM’s decoding algorithm.
Table 6. Defensive performance of BEAT on original Llama-3-8b-Instruct and PoisonedAlign variants under preference alignment (DPO) and supervised fine-tuning where malicious prompts are randomly sampled from Advbench (Zou et al., 2023) and MaliciousInstruct (Huang et al., 2024).
(a) Advbench
Model Metric Attack
Naive Attack Escape Characters Context Ignoring Fake Completion Combined Attack
DPO PoisonedAlign AUROC 0.71 0.77 0.91 0.95 0.99
TPR@FPR5% 0.16 0.24 0.58 0.73 0.98
SFT PoisonedAlign AUROC 0.35 0.41 0.43 0.71 0.85
TPR@FPR5% 0.00 0.01 0.00 0.05 0.25
Original
Llama-3-8b-Instruct
AUROC 0.60 0.64 0.58 0.79 0.93
TPR@FPR5% 0.19 0.19 0.10 0.33 0.81
(b) MaliciousInstruct
Model Metric Attack
Naive Attack Escape Characters Context Ignoring Fake Completion Combined Attack
DPO PoisonedAlign AUROC 0.66 0.75 0.96 0.95 0.99
TPR@FPR5% 0.12 0.21 0.82 0.75 1.00
SFT PoisonedAlign AUROC 0.39 0.48 0.50 0.77 0.87
TPR@FPR5% 0.00 0.01 0.00 0.04 0.18
Original
Llama-3-8b-Instruct
AUROC 0.66 0.71 0.67 0.84 0.92
TPR@FPR5% 0.13 0.19 0.16 0.45 0.68
5. Countermeasures
PoisonedAlign amplifies prompt injection vulnerabilities by poisoning the alignment process, making it a representative form of alignment-time data poisoning for backdoor attacks. We evaluate two recent and powerful defenses targeting backdoor or unalignment attacks on LLMs: BEAT (Yi et al., 2025) and BAIT (Shen et al., 2024), both of which are black-box detection methods designed to identify unaligned behavior or scan for compromised models. To assess their effectiveness against PoisonedAlign, we adopt the evaluation metrics defined in their original works. Our results show that while BEAT and BAIT provide partial mitigation, they either struggle under the dynamic and adaptive nature of prompt injection vulnerabilities enabled by PoisonedAlign, or exhibit limited effectiveness in distinguishing cleanly aligned models from those poisoned model by applying PoisonedAlign.
BEAT (Yi et al., 2025):  BEAT detects poisoned or triggered inputs at inference time by leveraging the probe concatenate effect. It appends each incoming input to a malicious probe and measures the change in the model’s refusal behavior. A substantial shift in the model’s response distribution, measured via metrics like Earth Mover’s Distance (EMD), signals the presence of a hidden trigger. To generate prompt injection samples, we employ separators listed in Table 1 to link probes with malicious prompts and these separators serve as backdoor mechanisms in this context. The underlying assumption is that successful prompt injections will cause notable distortions in the model’s output distribution. We consider two evaluation metrics: Area Under the Receiver Operating Characteristic Curve (AUROC) and True Positive Rate (TPR) at low False Positive Rate (FPR) for BEAT. AUROC shows the detector’s ability to separate triggered from non-triggered samples across thresholds. TPR at low FPR reflects how well the detector catches triggered samples while keeping false alarms low. Follow the default setting by BEAT, we control the FPR as 5%.
Table 6 presents the defensive performance of BEAT on both preference-aligned and supervised fine-tuned PoisonedAlign models based on Llama-3-8b-Instruct. Malicious prompts are drawn from Advbench (Zou et al., 2023) and MaliciousInstruct (Huang et al., 2024). We find that BEAT is effective when the prompt injection uses the separator from the Combined Attack, which serves as a strong ”backdoor” signal. For instance, when malicious prompts are sampled from MaliciousInstruct, BEAT achieves an AUROC of 0.99 and TPR@FPR5% of 1.00 on the preference alignment PoisonedAlign model. However, BEAT’s performance deteriorates significantly when other types of separators are used, especially on supervised fine-tuned models. For example, when using the Naive Attack’s separator with malicious prompts from Advbench, the AUROC drops to 0.35, and BEAT fails to detect any poisoned input on the supervised fine-tuned PoisonedAlign model.
BEAT’s effectiveness is also sensitive to the dataset from which malicious prompts are sampled. Under the same separator (e.g., context ignoring), BEAT reaches a TPR@FPR5% of 0.82 on supervised fine-tuned PoisonedAlign when prompts are from MaliciousInstruct, but this value drops to 0.58 when using prompts from Advbench. These results highlight BEAT’s limited robustness against PoisonedAlign. While PoisonedAlign enhances prompt injection vulnerabilities across diverse attack types, BEAT is overly specialized to the Combined Attack’s separator and struggles to generalize. Furthermore, BEAT introduces inference-time overhead for about 0.2 second, increasing the latency of each user query by approximately 1.2 times on two NVIDIA RTX A5000 GPUs.
BAIT (Shen et al., 2024):  BAIT operates at the model level, aiming to determine whether an LLM has been backdoored by inverting the backdoor target rather than the trigger. It systematically enumerates initial output tokens and tracks their causal dependencies through the model’s token generation process. A consistent sequence across diverse benign prompts indicates a potential backdoor. We use Q-SCORE (Shen et al., 2024) as the evaluation metric, which is a single numeric score between 0 and 1 that tells how strongly a candidate text sequence displays the “target-token causality” that only a backdoored LLM should reveal. We also keep all default settings for BAIT, including $k=5$ for top-K tokens to consider and $\phi=0.90$ as the threshold for Q-SCORE to generate binary prediction. We randomly sample 20 clean prompts from HH-RLHF (Bai et al., 2022) as a part of input of BAIT.
We apply BAIT to both clean and poisoned aligned Llama-3-8b-Instruct models, for both preference alignment and supervised finetuning. Table 7 reports the Q-SCOREs across different model variants. We observe that all models, including those fine-tuned on clean HH-RLHF, exhibit high Q-SCOREs exceeding the detection threshold. The differences between clean and poisoned models under the same alignment method are minimal, and supervised fine-tuning models tend to have slightly lower Q-SCOREs than their preference alignment counterparts. These results suggest that although BAIT can detect backdoor-like behavior in models fine-tuned with PoisonedAlign, it fails to effectively distinguish between clean and poisoned alignments. This limitation stems from the nature of PoisonedAlign, which does not implant a fixed target response but instead amplifies prompt injection vulnerabilities, making the model tend to complete the injected tasks.
Table 7. Q-SCORE of finetuned Llama-3-8b-Instruct models under preference alignment (DPO) and supervised fine-tuning, comparing outcomes achieved with clean finetuning versus PoisonedAlign. ‘Diff.’ denotes the difference between Q-SCOREs of poisoned alignment and clean alignment.
Alignment Method Clean Poisoned Diff.
DPO 0.9802 0.9798 -0.0004
SFT 0.9457 0.9458 0.0001
6. Discussion and Limitations
[…]